Weights. If the weights of 600 students are normally distributed with a mean of 50 kilograms and a variance of 16 kilograms
a. Determine the percentage of students with weights lower than 55 kilograms.
b. How many students have weights exceeding 52 kilograms?
1
Expert's answer
2020-12-21T17:29:14-0500

"\mu=50"

"\sigma=16"

"n=600"

a) "P(X<55)=P(Z<\frac{55-\mu}{\sigma})=P(Z<\frac{55-50}{16})=P(Z<0.3125)=0.6217"

wich is 62.17%

b) "P(X>52)=P(Z>\frac{52-50}{16})=P(Z>0.125)=1-P(Z<0.125)="

"=1-0.5478=0.4522"

"600\cdot0.4522=271"


Answer:

a) 62.17% of students with weights lower than 55 kilograms

b) 271 students have weights exceeding 52 kilograms